OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $16,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $3,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 1% simple annual interest and no down payment. (hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time (in years). Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Please expl

OpenStudy (anonymous):

oh hell no! that is not how one calculates payments when the interest is fixed.

OpenStudy (anonymous):

Payment Option 1: $16,000 – $3,000 = $13,000 after down payment monthly installment : $13,000 / 60months =$216.67/month Payment Option 2: interest : $16,000 * 0.01 * 6 = $960 total payment = $16000+$960=$16960 monthly installment=$16960/72months=$235.55/month Option 1 is lower